OpenStudy (anonymous):

Which pair of points are on the graph of y = -2x + 1? (2, 7) and (-1, 4) (3, -5) and (-4, -7) (6, 8) and (9, 3) (4, -7) and (-5,11) Find the midpoint of A(1, 4) and B(3, -6). Place your answer in (x,y) notation with NO spaces. (-1, -5) (2,-1) (4, -2) (2, -5)

OpenStudy (anonymous):

@Directrix

OpenStudy (aum):

y = -2x + 1 put the x and y values for each point in the above equation and and see if it is true. For example, the point (2,7). Put x = 2, y = 7 in y = -2x + 1 7 = -2 * 2 + 1 7 = - 4 + 1 = -3. This is NOT true. Therefore, (2,7) is NOT on the line y = -2x + 1.

OpenStudy (anonymous):

so is the first option the correct one?

OpenStudy (anonymous):

oh nvm

OpenStudy (aum):

I just showed you above that the point (2,7) is NOT on the line y = -2x + 1. So it is not the first choice. Try each point and see which pair of points lie on y = -2x + 1.

OpenStudy (anonymous):

ok

OpenStudy (aum):

For the second question, the midpoint of \((x_1, y_1)\) and \((x_2, y_2)\) is \(\large (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})\)